you have an equally likely chance of choosing an integer from 1 through 20 find the probability of choosing and odd number
Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

1/10
ODD NUMBERS FROM 1 THROUGH 20
1,3,5,7,9,11,13,15,17,19

Answer by Alan3354(69443)

you have an equally likely chance of choosing an integer from 1 through 20 find the probability of choosing and odd number
----------------
The statement is not clear, but there are 10 even numbers and 10 odd numbers.
choosing an odd # = 10/20 = 1/2